# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

### v3

@@ 6 Conclusion
In this work, we propose novel methodology for training a single model for amortized inference of SCMs. Cond-FiP not only generalizes to unseens in-distribution instances, but also to a wide range of OOD instances, including larger graphs, complex noise distributions, and real-world data. To the best of our knowledge, this is the first approach to demonstrate the feasibility of learning causal mechanisms in a reusable, foundational manner—paving the way for a paradigmatic shift towards the assimilation of causal knowledge across datasets.
Limitations. Our training is limited to synthetic additive noise SCMs due to the requirement of true noise variables for learning the dataset encoder. However, the conditional FiP decoder (see Section 3.2) does not rely on this assumption and can be applied to general SCMs given pretrained dataset embeddings. A promising direction for future work is to explore more general encoding schemes, such as self-supervised learning, or design an implicit in-context learning approach to remove the need for dataset embeddings via direct attention over the context (Mittal et al., 2024).
While Cond-FiP generalizes to larger graphs, it does not yet benefit from larger context sizes at inference (Section K.1), suggesting the need to scale both the model and training data for richer contexts. Additionally, although Cond-FiP performs well on generating interventional samples, it doesn’t perform well on counterfactual generation (Section K.2). Future work will explore scaling Cond-FiP to larger problem instances and application for more complex tasks (counterfactual generation) in real-world scenarios.
@@ Acknowledgements
We thanks the members of the Machine Intelligence team at Microsoft Research for helpful discussions. We also thank Sarthak Mittal for their suggestion to benchmark Cond-FiP in the scarce data regime. Further, we thank Moksh Jain for their feedback on the draft. Part of the experiments were enabled by the Digital Research Alliance of Canada (https://alliancecan.ca/en) and Mila cluster (https://docs.mila.quebec/index.html). Divyat Mahajan acknowledges support via FRQNT doctoral scholarship (https://doi.org/10.69777/354785) for his graduate studies.
Impact Statement
@@ A.1 DAG-Attention Mechanism
In FiP [Scetbon et al., 2024] the authors propose to leverage the transformer architecture to learn SCMs from observations. By reparameterizing an SCM according to a topological ordering induced by its graph, the authors show that any SCM can be reformulated as a fixed-point problem of the form ${\bm{X}}=\bm{H}({\bm{X}},{\bm{N}})$ where $\bm{H}$ admits a simple triangular structure:
@@ A.3 Inference with Cond-FiP
In order to intervene on this predicted SCM, we simply have to modify in place the predicted function. For example, assume that we want to perform the following intervention $\text{do}(X_{i})=a$. Then, to obtain the intervened SCM, we define a new function $z\to\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})$ defined for any $z\in\mathbb{R}^{d}$ as: $[\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})]_{j}:=[\mathcal{T}(z,D_{\bm{X}},\mathcal{G})]_{j}$ if $j\neq i$ and $[\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})]_{i}:=a$.
@@
Appendix B Details on Experiment Setup with AVICI Benchmark
@@ B.1 AVICI Benchmark
We use the synthetic data generation procedure proposed by Lorch et al. [2022] to generate SCMs in our empirical study. It provides access to a wide variety of SCMs, hence making it an excellent setting for amortized training.
  • Graphs: We have the option to sample graphs as per the following schemes: Erods-Renyi [Erdos and Renyi, 1959], scale-free models [Barabási and Albert, 1999], Watts-Strogatz [Watts and Strogatz, 1998], and stochastic block models [Holland et al., 1983].
@@
We provide further details on the shift in the support of parameters for functional mechanisms below. For complete details please refer to Table 3, Appendix in Lorch et al. [2022].
@@
B.2 Model Architecture and Training Details
For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. The models were trained for a total of $10k$ epochs with the Adam optimizer [Paszke et al., 2017], where we used a learning rate of $41e-$ 4 and a weight decay of $95e-$ 9. Each epoch contains $\simeq 400$ randomly generated datasets from the distribution ${\mathbb{P}}_{\text{IN}}$. We also use the EMA implementation of [Karras et al., 2023] to train our models.
@@ B.2 Model Architecture and Training Details
  • Each batch consists of $n=400$ samples with dimension $d=20$ requiring less than $1$ MiB of data in FP32 precision.
  • Storing the model on the GPU requires under 100 MiB.
@@
Compared to the baselines, Cond-FiP has similar memory requirements to DECI [Geffner et al., 2022] and FiP [Scetbon et al., 2024], as all three train neural networks of comparable size. The main exception is DoWhy [Blöbaum et al., 2022], which fits simpler models for each node, but this approach does not scale well as the graph size increases.
Computational Cost.
B.3 Code Repository
We plan to open-source the code along with comprehensive documentation to facilitate reproducibility of our experiments. For the submission, we have prepared an anonymized version of the codebase, which can be accessed via this link: https://anonymous.4open.science/r/neurips_2025_condfip-1277/.
Please note that while the codebase is not directly executable, it provides full access to the implementation of all components of our framework:
  • cond-fip/models contains the implementation of the transformer-based encoder and the Cond-FIP architecture.
  • cond-fip/tasks includes the training and inference methods associated with our framework.
@@ Appendix D Experiments on Sensitivity to Distribution Shifts on AVICI benchmark
In Appendix C (Table 1Table 2Table 3), we tested OOD genrealization with datasets sampled from SCM following a different distribution (LIN OUT, RFF OUT) than the datasets used for training Cond-FiP (LIN IN, RFF IN). We now analyze how sensitive is Cond-FiP to distribution shifts by comparing its performance across scenarios as the severity of the distribution shift is increased.
@@
Tables 45, and 6 provide results for the case of controlling shift via causal mechanisms, for the task of noise prediction, sample generation, and interventional generation respectively. We find that the performance of Cond-FiP does not change much as we increase $\alpha$, indicating that Cond-FiP is robust to the varying levels of distribution shits in causal mechanisms.
However, for the case of controlling shift via noise variables (Table 78, and 9) we find that Cond-FiP is quite sensitive to the varying levels of distribution shift in noise variables. The performance of Cond-FiP degrades with increasing magnitude of the shift ($\alpha$) for all the tasks.
@@
Table 4: Results for Noise Prediction under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 5: Results for Sample Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 6: Results for Interventional Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 7: Results for Noise Prediction under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
@@
Table 8: Results for Sample Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
@@
Table 9: Results for Interventional Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.

### v4

@@ 5 Conclusion
In this work, we propose novel methodology for training a single model for amortized inference of causal mechanisms in SCMs. Cond-FiP not only generalizes to unseen in-distribution instances, but also to a wide range of OOD instances, including larger graphs, unknown causal graphs, complex noise distributions, and real-world data. To the best of our knowledge, this is the first approach to demonstrate the feasibility of learning causal mechanisms in a reusable, foundational manner—paving the way for a paradigmatic shift towards the assimilation of causal knowledge across datasets.
Limitations. Our training is limited to synthetic additive noise SCMs due to the requirement of true noise variables for learning the dataset encoder. However, the conditional FiP decoder (see Section˜3.2) does not rely on this assumption and can be applied to general SCMs given pretrained dataset embeddings. A promising direction for future work is to explore more general encoding schemes, such as self-supervised learning, or design an implicit in-context learning approach to remove the need for dataset embeddings via direct attention over the context (Mittal et al., 2024). We believe our framework can serve as a good motivation for future works that incorporate real-world datasets during training as well.
While Cond-FiP generalizes to larger graphs, it does not yet benefit from larger context sizes at inference (Section˜I.1), suggesting the need to scale both the model and training data for richer contexts. Additionally, although Cond-FiP performs well on generating interventional samples, it doesn’t perform well on counterfactual generation (Section˜I.2). Future work will explore scaling Cond-FiP to larger problem instances and application for more complex tasks (counterfactual generation) in real-world scenarios.
@@ Acknowledgements
We thank the members of the Machine Intelligence team at Microsoft Research for helpful discussions. We also thank Sarthak Mittal for their suggestion to benchmark Cond-FiP in the scarce data regime. Further, we thank Moksh Jain for their feedback on the draft. Part of the experiments were enabled by the Digital Research Alliance of Canada (https://alliancecan.ca/en) and Mila cluster (https://docs.mila.quebec/index.html). Divyat Mahajan acknowledges support via FRQNT doctoral scholarship (https://doi.org/10.69777/354785) for his graduate studies.
Broader Impact Statement
@@ A.1 DAG-Attention Mechanism
In FiP (Scetbon et al., 2024) the authors propose to leverage the transformer architecture to learn SCMs from observations. By reparameterizing an SCM according to a topological ordering induced by its graph, the authors show that any SCM can be reformulated as a fixed-point problem of the form ${\bm{X}}=\bm{H}({\bm{X}},{\bm{N}})$ where $\bm{H}$ admits a simple triangular structure:
@@ A.3 Inference with Cond-FiP
In order to intervene on this predicted SCM, we simply have to modify in place the predicted function. For example, assume that we want to perform the following intervention $\text{do}(X_{i})=a$. Then, to obtain the intervened SCM, we define a new function $z\to\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})$ defined for any $z\in\mathbb{R}^{d}$ as: $\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})]_{j}:=[\mathcal{T}(z,D_{\bm{X}},\mathcal{G})]_{j}$ if $j\neq i$ and $\mathcal{T}^{\text{do}(X_{i})=a}(z,D_{\bm{X}},\mathcal{G})]_{i}:=a$.
@@
A.4 Pseduo Code
Algorithm 1 Cond-FiP Part 1: Dataset Encoder $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$
1:Observational dataset $D_{\bm{X}}\in\mathbb{R}^{n\times d}$, causal graph ${\bm{\mathcal{G}}}\in\{0,1\}^{d\times d}$
2:Dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$
3:Linear Embedding: Apply a learned linear map $L:\mathbb{R}^{1}\to\mathbb{R}^{d_{h}}$ to each sample:
$L(D_{\bm{X}})=[\bm{X}_{1}W_{L},\ldots,\bm{X}_{n}W_{L}]^{\top}\in\mathbb{R}^{n\times d\times d_{h}}.$
4:Alternating Attention: Pass $L(D_{\bm{X}})$ through a stack of transformer blocks that alternate between:
  • Sample-wise attention (unmasked) across the $n$ observations, capturing global context.
  • Node-wise attention with mask defined as per the causal graph.
    $M_{ij}=\begin{cases}0,&\text{if }G_{ij}=1,\\ -\infty,&\text{otherwise}.\end{cases}$
5:Contextual Representation: After alternating attention blocks, obtain dataset embeddings
$\mu(D_{\bm{X}},{\bm{\mathcal{G}}})=E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}.$
6:Prediction Head: Apply a two-layer MLP $H$ to map embeddings to estimated noise or function values:
$\hat{Y}=H(\mu(D_{\bm{X}},{\bm{\mathcal{G}}}))\in\mathbb{R}^{n\times d}.$
7:Training Objective: Under the additive noise model assumption, minimize the MSE loss:
$\mathcal{L}_{E}=\mathbb{E}_{S\sim P_{S}}\!\left[\|\bm{F}(D_{\bm{X}})-\hat{Y}\|_{2}^{2}\right].$
8:Output: Return the dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$
Algorithm 2 Cond-FiP Part 2: Conditional Fixed-Point Decoder ${\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$
1:Input point $\bm{z}\in\mathbb{R}^{d}$, dataset $D_{\bm{X}}\in\mathbb{R}^{n\times d}$, causal graph ${\bm{\mathcal{G}}}\in\{0,1\}^{d\times d}$, encoder output $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$
2:Learned causal model: ${\mathcal{T}}(z,D_{\bm{X}},{\bm{\mathcal{G}}})$
3:Pooling: Aggregate the encoder representations across samples:
$\mu^{\prime}(D_{\bm{X}},{\bm{\mathcal{G}}})=\text{MaxPool}_{n}\big(E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\big)\in\mathbb{R}^{d\times d_{h}}.$
4:Context-Specific Codebooks: Compute dataset-conditioned codebooks and positional embeddings:
$C(D_{\bm{X}},{\bm{\mathcal{G}}})=\mu^{\prime}(D_{\bm{X}},{\bm{\mathcal{G}}})W_{C},\quad P(D_{\bm{X}},{\bm{\mathcal{G}}})=\mu^{\prime}(D_{\bm{X}},{\bm{\mathcal{G}}})W_{P}.$
5:Input Embedding: Embed the input $\bm{z}$ via element-wise modulation:
$\bm{z}_{\text{emb}}=[\,z_{1}C_{1}(D_{\bm{X}},{\bm{\mathcal{G}}}),\ldots,z_{d}C_{d}(D_{\bm{X}},{\bm{\mathcal{G}}})\,]+P(D_{\bm{X}},{\bm{\mathcal{G}}}).$
6:Initialization: Set initial latent state (which will be transformed into the estimated causal variable) using the dataset embedding:
$n_{0}=\mu^{\prime}(D_{\bm{X}},{\bm{\mathcal{G}}})W_{n_{0}}.$
7:Fixed-Point Iterations: For $\ell=0,\ldots,L-1$, apply DAG-Attention and adaptive normalization:
$\bm{n}_{\ell+1}=h\!\left(\text{DAM}(\bm{n}_{\ell},\bm{z}_{\text{emb}})\,\bm{z}_{\text{emb}}+\bm{n}_{\ell}\right),$
where $h(\cdot)$ includes adaptive LayerNorm whose scale and shift depend on $\mu^{\prime}(D_{\bm{X}},{\bm{\mathcal{G}}})$.
8:Output Projection: After $L$ iterations, project the final state:
${\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})=\bm{n}_{L}W_{\text{out}}\in\mathbb{R}^{d}.$
9:Training Objective: Minimize reconstruction loss over random samples $z\sim P_{X}$:
$\mathcal{L}_{T}=\mathbb{E}_{S\sim P_{S},\,\bm{z}\sim P_{X}}\!\left[\|\,{\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})-\bm{F}(\bm{z})\,\|_{2}^{2}\right].$
Appendix B Details on Experiment Setup
@@ B.1 AVICI Benchmark
We use the synthetic data generation procedure proposed by Lorch et al. (2022) to generate SCMs in our empirical study. It provides access to a wide variety of SCMs, hence making it an excellent setting for amortized training.
  • Graphs: We have the option to sample graphs as per the following schemes: Erods-Renyi (Erdos & Renyi, 1959), scale-free models (Barabási & Albert, 1999), Watts-Strogatz (Watts & Strogatz, 1998), and stochastic block models (Holland et al., 1983).
@@
We provide further details on the shift in the support of parameters for functional mechanisms below. For complete details please refer to Table 3, Appendix in Lorch et al. (2022).
@@
B.2 CSuite Benchmark
CSuite (Geffner et al., 2022) is a collection of synthetic structural causal models (SCMs) designed to evaluate causal discovery and effect estimation methods. The benchmark covers a diverse set of settings by varying graph structures, functional forms, and noise distributions, thereby testing models under both linear and nonlinear, Gaussian and non-Gaussian conditions, etc. We use the lingauss, linexp, nonlingauss, nonlin-simpson, symprod-simpson, large-backdoor, and weak-arrow tasks from their paper.
  • Noise is sampled with equal probability from either N(2,1)N(-2,1) and $N(2,1)$.
  • Noise is sampled with equal probability from either N(2,2)N(-2,2) and $N(2,2)$.
  • Noise is sampled with equal probability from either N(2,1)N(-2,1) and $N(2,2)$.
  • Noise is sampled with equal probability from either N(5,1)N(-5,1) and $N(5,1)$.
  • Noise is sampled with equal probability from either N(5,2)N(-5,2) and $N(5,2)$.
  • Noise is sampled with equal probability from either N(5,1)N(-5,1) and $N(5,2)$.
This leads to a total of $12$ experimental setting with $6$ different GMM noise distribution for both the Large Backdoor and Weak Arrow datasets from the CSuite benchmark.
B.3 Model Architecture and Training Details
For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. The models were trained for a total of $10k$ epochs with the Adam optimizer (Paszke et al., 2017), where we used a learning rate of 1e41e-4 and a weight decay of 5e95e-9. Each epoch contains $\simeq 400$ randomly generated datasets from the distribution ${\mathbb{P}}_{\text{IN}}$. We also use the EMA implementation of (Karras et al., 2023) to train our models.
@@ B.3 Model Architecture and Training Details
  • Each batch consists of $n=400$ samples with dimension $d=20$ requiring less than $1$ MiB of data in FP32 precision. Also, storing the model on the GPU requires under 100 MiB.
@@
Compared to the baselines, Cond-FiP has similar memory requirements to DECI (Geffner et al., 2022) and FiP (Scetbon et al., 2024), as all three train neural networks of comparable size. The main exception is DoWhy (Blöbaum et al., 2022), which fits simpler models for each node, but this approach does not scale well as the graph size increases.
@@ Appendix D Experiments on Sensitivity to Distribution Shifts on AVICI benchmark
In Appendix˜C (Table˜3Table˜4Table˜5), we tested OOD genrealization with datasets sampled from SCM following a different distribution (LIN OUT, RFF OUT) than the datasets used for training Cond-FiP (LIN IN, RFF IN). We now analyze how sensitive is Cond-FiP to distribution shifts by comparing its performance across scenarios as the severity of the distribution shift is increased.
@@
Tables 67, and 8 provide results for the case of controlling shift via causal mechanisms, for the task of noise prediction, sample generation, and interventional generation respectively. We find that the performance of Cond-FiP does not change much as we increase $\alpha$, indicating that Cond-FiP is robust to the varying levels of distribution shits in causal mechanisms.
However, for the case of controlling shift via noise variables (Table 910, and 11) we find that Cond-FiP is quite sensitive to the varying levels of distribution shift in noise variables. The performance of Cond-FiP degrades with increasing magnitude of the shift ($\alpha$) for all the tasks.
@@
Table 6: Results for Noise Prediction under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 7: Results for Sample Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜4. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 8: Results for Interventional Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜5. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
@@
Table 9: Results for Noise Prediction under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
@@
Table 10: Results for Sample Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜4. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
@@
Table 11: Results for Interventional Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table˜5. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.